OpenStudy (anonymous):

an angle measures 46 degrees what is the measure of its complement. an angle measures 65 degrees. what is the measure of its supplement.

OpenStudy (anonymous):

can someone help me please

OpenStudy (anonymous):

complement is 90 degree 46+x=90 x= 90-46 x =44 supplement is 180 dgree 65+x=180 x =180-65 x= 115